OpenStudy (anonymous):

Use Cramer’s rule to solve the following system. x-y+2z=3 2x+3y+z=9 -x-y+3z=11

OpenStudy (anonymous):

Find Dx and Dy and Dz

OpenStudy (anonymous):

3 -1 2 9 3 1 11 -1 3 Find determinant of this matrix that is your Dx

OpenStudy (anonymous):

any question?

OpenStudy (anonymous):

Dx/D is x Likewise for y,z

OpenStudy (anonymous):

Do you want me go through the problem?

OpenStudy (anonymous):

Yes please

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

3 -1 2 9 3 1 11 -1 3 Let's Find the Determinant of this matrix, Dx

OpenStudy (anonymous):

Do you know how to find determinant?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

is the determinant -38

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Now Find Dy

OpenStudy (anonymous):

57

OpenStudy (anonymous):

Dz

OpenStudy (anonymous):

76

OpenStudy (anonymous):

then find determinant of D 1 1 2 2 3 1 1 1 3

OpenStudy (anonymous):

x= Dx/D y= Dy/D z=Dz/D

OpenStudy (anonymous):

wouldn't it be 1 -1 2 2 3 1 1 1 3 because the first system is x-y+2z=3

OpenStudy (anonymous):

Yes, I missed -,sry

OpenStudy (anonymous):

solution set is x=-38,y=57,z=76